OpenStudy (anonymous):

Which expression represents the area of the rectangle? one side is 2/3 and the other side

OpenStudy (anonymous):

i need help i need to be guided toward the right way

OpenStudy (anonymous):

The question is unclear.

OpenStudy (anonymous):

http://www.mathsisfun.com/area.html

OpenStudy (anonymous):

the choices r F. 12n - 18 G. 4n - 6 H. 6n -9 I. 4n - 4

OpenStudy (anonymous):

That is not correct. The area of a rectangle is width x height. Nothing regarding n is even mentioned.

OpenStudy (anonymous):

well thats what it has but i told you one of the sides of the rectangle is (6n-9) and the other 2/3

OpenStudy (anonymous):

You made no mention of that. You wrote... Which expression represents the area of the rectangle? one side is 2/3 and the other side

OpenStudy (anonymous):

oh i am so sorry i meant to write that sorry i am doing other things so i did not even realize

OpenStudy (anonymous):

There is no such thing as multitasking. Just switching your attention rapidly. The more switching the less focus that can be given to all task.

OpenStudy (anonymous):

A = L x H A = 2/3(6n - 9) divide by 3 to get rid of fraction 3A = 2(6n - 9) 3A = 12n - 18 A = (12n - 18)/3 A = 4n - 6